OpenStudy (anonymous):

2x2 - 5x – 10 = 0 solve using the quadratic formula.

OpenStudy (anonymous):

the formula is \[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

ur a here is 2; b=-5 and c=10 plug in and solve

OpenStudy (anonymous):

c=-10*

OpenStudy (anonymous):

2x^2-5x-10=0 Quadratic Formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] a=2 b=-5 c=-10 \[x=\frac{5\pm\sqrt{25+80}}{4}=\frac{5\pm\sqrt{105}}{4}\]

OpenStudy (anonymous):

\[2x^2 - 5x - 10 =0 \] We know the quadratic formula is of the form... \[x = \frac{-b \pm \sqrt{(b^2-4ac)}}{2a} \] Let's try to determine what these variables correspond to in our given equation. \[ a = 2\\b = -5\\c = -10 \] So, now we can plug and solve. \[x = \frac{5 \pm \sqrt{(25+80)}}{4} \]